# PBR-NeRF: Inverse Rendering with Physics-Based Neural Fields

### v1

@@ 1 Introduction
Figure 1: Improving inverse rendering with physics-based priors. The proposed PBR-NeRF significantly outperforms the NeILF++ [51] baseline by simply using our novel Conservation of Energy and NDF-weighted Specular Losses. Our physics-based losses correct “baked-in” specular highlights misrepresented in the diffuse lobe (highlighted areas) by (1) enforcing energy conservation and (2) accurately separating specular and diffuse reflections. The result is a more realistic, physically consistent material and lighting estimation.
@@
We achieve these goals through two novel physics-based losses applied directly to the Disney BRDF. Our first loss enforces energy conservation by penalizing physically invalid BRDFs that reflect more energy than received. Our second loss promotes the separation of specular and diffuse BRDF lobes, which is crucial for accurately modeling highly specular surfaces. Without constraints, the diffuse lobe (which defines the albedo) tends to overcompensate near symmetric reflection angles, where both lobes contribute to the aggregate BRDF. By penalizing the diffuse lobe’s magnitude, weighted by the normal distribution function to target specular angles, this loss indirectly encourages the specular lobe to capture and fully “explain” specular highlights effectively.
We thoroughly validate PBR-NeRF on two major benchmarks for inverse rendering and novel view synthesis: the NeILF++ [51] and DTU [15] datasets. Our experiments show that PRB-NeRF sets the new state of the art for material estimation in inverse rendering, while maintaining or surpassing the competitive novel view synthesis performance of its baseline. Crucially, our model delivers faithful albedo estimates that remain consistent for a given scene across diverse illuminations, thanks to its effective diffuse-specular decomposition. This decomposition also improves metallicness and roughness estimates as an additional benefit, further highlighting the effectiveness of our approach.
@@ 2 Related Work
Neural fields for novel view synthesis  have achieved impressive progress by representing 3D scenes as continuous volumetric fields [23]. This family of methods achieves photorealistic synthesis from unseen viewpoints through differentiable volume rendering, which estimates outgoing radiance by alpha-compositing color over sampled points along each camera ray. We concisely review representative works here and refer the reader to [41] for a comprehensive overview of the associated literature. Scene geometry is commonly represented using density fields [23, 52, 4, 5, 24, 6, 42, 28], 3D Gaussians [18, 12, 39], or signed distance fields (SDFs) [46, 45, 36, 49, 50] for images and other optical modalities [7, 14, 30]. While the density fields in NeRF and 3D Gaussians can model arbitrary complex scene geometries, they fundamentally lack well-defined surfaces and normals. These particle-centric representations result in ambiguous surface definitions, which hinder their applicability to PBR. In contrast, SDFs explicitly define surfaces as the zero-level set of a neural network, providing accurate surface and normal estimation. This surface-centric representation makes SDFs inherently more compatible with inverse rendering tasks which rely on PBR and require precise surface definitions.
@@
Figure 2: Overview of our PBR-NeRF architecture for neural inverse rendering. Our two novel physics-based material losses, i.e. $\mathcal{L}_{\text{cons}}$ for conservation of energy and $\mathcal{L}_{\text{spec}}$ for disentangling the diffuse and specular BRDF components, are derived in Sec. 3.2. The complete PBR-NeRF model comprises multiple neural fields, which are optimized in a stage-wise fashion: a standard NeRF+SDF modeling geometry and radiance, a neural incident light field (NeILF) modeling spatially-varying illumination, and a BRDF field modeling materials via the Disney BRDF [11]. Our physics-based losses provide the BRDF field with valuable inductive biases for improved material estimation, which help resolve to a large extent the inherent material-lighting ambiguity in inverse rendering and thus benefit the incident light field as well. Consequently, novel scene views synthesized with our model enjoy state-of-the-art quality.
@@
Neural inverse rendering  predates neural fields [33, 19, 37, 48, 31, 22] and learns to predict 3D geometry, materials, and illumination from 2D images. Decomposed neural fields [35, 38, 47, 27] similarly extend conventional neural fields to estimate physical scene properties such as materials and illumination, through inverse rendering or intrinsic decomposition. Compared to NeRF and 3DGS, the above additional material and lighting estimation enables complex downstream tasks such as relighting [43, 29, 21], material editing, and appearance manipulation. The decomposition of reflected radiance into materials and illumination also helps to address issues such as “baked-in” highlights in predicted colors by properly modeling view-dependent effects. However, inverse rendering introduces significant complexity due to the material-lighting-geometry ambiguity, which is harder to lift than the geometry-radiance ambiguity of standard novel view synthesis. Many approaches simplify this task by assuming known lighting, environment map lighting [53, 54, 25, 8, 9], known geometry [44], or basic Lambertian material models. State-of-the-art inverse rendering approaches jointly model (1) direct and indirect lighting from near and far sources [44, 51, 40], and (2) spatially-varying microfacet BRDFs [34, 44, 51]. For computational efficiency, techniques such as light caching [55, 44, 20, 51, 16, 3] and incident light sampling [44, 51, 3] are crucial to avoid costly recursive light tracing.
@@
NeILF [44] proposed joint incident light and BRDF fields, improving on environment-map-based approaches such as NeRV [34] by handling near-field lights and mixed lighting settings in addition to indirect lighting and occlusions. Building on this, NeILF++ [51] introduced inter-reflectable light fields, combining the NeILF and BRDF fields with an SDF for geometry estimation. This approach uses the SDF’s volume-rendered outgoing radiance to supervise the incident radiance predicted by the NeILF. The reflection loss of [51] harnesses the SDF’s strong novel view synthesis performance to steer the ill-posed inverse rendering, achieving state-of-the-art material estimation among neural inverse rendering works. While largely effective, current inverse rendering methods including [44, 51] lack informative material priors for disambiguating specular from diffuse properties of scene surfaces. This typically results in incorrectly attributing specular highlights to alleged changes in diffuse albedo, even though such errors do not affect view synthesis results.
@@
Material priors for inverse rendering  are crucial for constraining material estimation and reducing material-lighting ambiguities. Basic BRDF smoothness priors are used in [44, 51, 16, 3] to discourage abrupt spatial changes in material properties. NeILF [44] and NeILF++ [51] also use a Lambertian prior for regularizing the Disney BRDF’s metallicness and roughness parameters. However, the strong assumption of perfectly rough, non-metallic materials induced by this Lambertian prior limits BRDF expressiveness. For instance, specular effects often appear in the predicted BRDF’s diffuse lobe due to the Lambertian prior suppressing the specular lobe. We address these limitations by proposing two novel physics-based losses that enhance the physical validity and accuracy of estimated BRDFs without sacrificing expressiveness.
@@ 3.2 Physics-Based Losses
NDF-weighted Specular Loss.  Our second physics-based loss targets the imbalance between the diffuse and specular lobes frequently observed in inverse rendering methods, such as NeILF++ [51], which often assume Lambertian reflection. Real-world materials, however, often violate idealized Lambertian behavior, resulting in “baked-in” specular highlights where the diffuse lobe $f_{d}$ compensates for insufficient specular reflection. This imbalance causes incorrect diffuse behavior at non-specular angles, degrading material estimation quality.
To mitigate this issue, we propose the NDF-weighted Specular Loss to penalize excessive diffuse reflection in specular regions. When specular highlights are “baked-in”, the aggregate BRDF value $f_{r}$ may be accurate for specular directions, but the oversized diffuse lobe $f_{d}$ overcompensates the specular lobe $f_{s}$. We define our loss as
$\mathcal{L}_{\text{spec}}=\\ \frac{1}{|S_{L}|}\sum_{\boldsymbol{\omega}_{i}\in S_{L}}\operatorname{softmax}\left(\frac{\operatorname{detach}\left(D\left(\frac{\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}}{||\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}||}\right)\right)}{T_{\text{spec}}}\right)f_{d},$ (8)
@@
Our NDF-weighted Specular Loss complements the standard RGB rendering loss by pushing the specular lobe $f_{s}$ to compensate for the penalized diffuse lobe $f_{d}$, creating a stronger separation between diffuse and specular reflections. When the specular lobe is properly recovered, the RGB rendering loss prevents the diffuse lobe from vanishing. We note that our loss is meant to fine-tune material estimation after initial convergence with the RGB rendering loss. A proper weighting of the opposing RGB rendering loss and other regularization terms ensures that the diffuse lobe is not arbitrarily suppressed by our specular loss. As with $\mathcal{L}_{\text{cons}}$, the specular loss also benefits lighting estimation. An oversized diffuse lobe can lead to lower estimates of light intensities, while excessive diffuse reflection forces the estimated incident light to produce complex, view-dependent specular effects. Together, these losses improve the accuracy of material and lighting estimation significantly, resulting in more realistic scene reconstructions, as we detail in Sec. 4.
@@ 3.3 Joint Optimization
Following NeILF++ [51], we optimize the scene reconstruction in three phases. We briefly summarize the three phases here (See Appendix Sec. C.2 for full details):
@@
$\mathcal{L}_{\text{mat}}=\lambda_{\text{pbr}}\mathcal{L}_{\text{pbr}}+\mathcal{L}_{\text{ref}}+\lambda_{\text{smth}}\mathcal{L}_{\text{smth}}+\mathcal{L}_{\text{mat,physics}},$ (9)
@@ 4 Experiments
Table 1: Comparison against state-of-the-art methods on the NeILF++ dataset [51]. Performance in material estimation and novel view synthesis is compared across different scenes (City, Studio, and Castel) and types of illumination (Env: global environment map, Mix: mixed lighting with environment map, point and area light sources) using the PSNR metric. †: our reproduced results.
Predicted Quantity Method Env Mix Mean
City Studio Castel City Studio Castel
RGB PhySG [53] 24.82 25.65 27.24 24.38 24.04 25.81 25.32
SG-ENV [53] 31.01 29.46 32.34 27.20 25.88 27.70 28.93
NeILF [44] 33.77 31.07 35.28 31.11 28.59 32.11 31.99
NeILF++ [51] 31.53 29.77 33.87 29.64 27.50 30.74 30.51
PBR-NeRF (ours) 32.05 30.67 34.58 29.13 27.13 30.03 30.60
Roughness PhySG [53] 6.62 11.29 6.22 6.27 6.83 6.14 7.23
SG-ENV [53] 9.61 17.64 9.74 8.77 12.58 9.14 11.25
NeILF [44] 17.48 18.30 13.40 17.05 16.27 16.44 16.49
NeILF++ [51] 21.22 22.01 19.91 22.19 23.43 22.20 21.83
PBR-NeRF (ours) 22.20 23.19 22.49 22.19 22.30 22.36 22.46
Metallicness PhySG [53] 8.72 7.97 8.35 8.67 8.95 8.76 8.57
SG-ENV [53] 17.01 16.40 16.39 15.44 14.25 14.49 15.66
NeILF [44] 18.22 19.11 10.29 18.42 18.43 17.34 16.98
NeILF++ [51] 19.49 18.17 17.41 20.80 17.98 22.20 19.34
PBR-NeRF (ours) 21.94 21.12 21.24 21.63 17.40 20.50 20.64
Albedo PhySG [53] 15.01 16.96 16.13 14.16 12.43 14.29 14.83
SG-ENV [53] 22.38 20.74 22.21 16.92 13.16 16.69 18.68
NeILF [44] 15.59 15.48 12.95 17.39 16.88 17.37 15.94
NeILF++ [51] 18.21 19.50 17.73 16.37 14.49 17.43 17.29
PBR-NeRF (ours) 19.42 19.15 19.50 19.33 18.36 18.78 19.09
@@ 4.1 Experimental Setup
Implementation Details.  We implement PBR-NeRF on top of NeILF++ [51] using PyTorch [2] and keep all hyperparameters from NeILF++ the same, except for $|S_{L}|=256$. We train for 30K iterations with an initial learning rate of 0.002, reducing it by a factor of 5 every 10K iterations. Training runs on a single NVIDIA RTX 4090, with runtimes ranging from 1 to 3 hours. Using grid search, we identify two sets of weights for our physics-based losses: $\lambda_{\text{cons}}=0.01$ and $\lambda_{\text{spec}}=0.5$ for the NeILF++ dataset, and $\lambda_{\text{cons}}=0.01$ and $\lambda_{\text{spec}}=0.01$ for DTU. More hyperparameter details are provided in Appendix Sec. C.3.
Datasets.  We evaluate our method on the NeILF++ [51] and DTU [15] datasets. We use the NeILF++ synthetic dataset to evaluate our material-illumination estimation, assuming known geometry. This dataset provides ground-truth geometry, RGB, albedo, metallicness, and roughness for a scene rendered under 6 static illumination patterns, including three environment maps (Env) and three mixed-lighting scenarios with additional point and area light sources (Mix). Each pattern includes 96 RGB HDR $1600\times 1200$ images, split into 87 training and 9 validation images. Ground-truth geometry and poses are used in training, while ground-truth color, albedo, metallicness, and roughness are only used for evaluation. DTU captures 15 real-world objects in a controlled lab setting. Each scene includes 49 RGB LDR $1600\times 1200$ images, split into 44 training and 5 validation images. We only qualitatively evaluate material estimation on DTU because it contains LDR images and does not provide ground-truth materials.
@@
Table 2: Comparison for novel view synthesis against state-of-the-art methods on DTU [15]. Performance is compared across different scenes using the PSNR metric. †: our reproduced results. $\ddagger$: mean result over 3 runs.
Method \ DTU scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
RGB PSNR NeRFactor [54] 23.24 21.91 23.33 26.86 22.70 24.71 27.59 22.56 20.45 25.08 26.30 25.14 21.35 26.44 26.53 24.28
PhySG [53] 17.38 15.11 20.65 18.71 18.89 18.47 18.08 21.98 17.31 20.67 18.75 17.55 21.20 18.78 23.16 19.11
Neural-PIL [9] 20.67 19.51 19.12 21.01 23.70 18.94 17.05 20.54 19.88 19.67 18.20 17.75 21.38 21.69 - 19.94
NeILF++ [51$\ddagger$ 25.85 21.93 25.42 27.99 29.76 27.43 26.54 28.37 23.71 28.28 31.56 30.45 26.86 31.32 32.19 27.85
PBR-NeRF (ours)$\ddagger$ 25.97 21.91 27.32 27.95 29.88 27.45 26.96 30.88 24.49 27.53 31.71 30.53 26.91 30.55 33.22 28.22

### v2

@@ 1 Introduction
Figure 1: Improving inverse rendering with physics-based priors. The proposed PBR-NeRF significantly outperforms the NeILF++ [51] baseline simply by using our novel Conservation of Energy and NDF-weighted Specular Losses. Our physics-based losses correct “baked-in” specular highlights misrepresented in the diffuse lobe (highlighted areas) by (1) enforcing energy conservation and (2) accurately separating specular and diffuse reflections. The result is a more realistic, physically consistent material and lighting estimation.
@@
In particular, we achieve these goals through two novel physics-based losses applied directly to the Disney BRDF. Our first loss enforces energy conservation by penalizing physically invalid BRDFs that reflect more energy than received. Our second loss promotes the separation of specular and diffuse BRDF lobes, which is crucial for accurately modeling highly specular surfaces. Without constraints, the diffuse lobe (which defines the albedo) tends to overcompensate near symmetric reflection angles, where both lobes contribute to the aggregate BRDF. By penalizing the diffuse lobe’s magnitude, weighted by the normal distribution function to target specular angles, this loss indirectly encourages the specular lobe to expand and fully “explain” specular highlights effectively.
We thoroughly validate PBR-NeRF on two major benchmarks for inverse rendering and novel view synthesis: the NeILF++ [51] and DTU [15] datasets. Our experiments show that PBR-NeRF sets the new state of the art for material estimation in inverse rendering, while maintaining or surpassing the competitive novel view synthesis performance of its baseline. Crucially, our model delivers faithful albedo estimates that remain consistent for a given scene across diverse illuminations, thanks to its effective diffuse-specular decomposition. This decomposition also improves metallicness and roughness estimates as an additional benefit, further highlighting the effectiveness of our approach.
@@ 2 Related Work
Neural fields for novel view synthesis  have achieved impressive progress by representing 3D scenes as continuous volumetric fields [23]. This family of methods achieves photorealistic synthesis from unseen viewpoints through differentiable volume rendering, which estimates outgoing radiance by alpha-compositing color over sampled points along each camera ray. We concisely review representative works here and refer the reader to [41] for a comprehensive overview of the associated literature. Scene geometry is commonly represented using density fields [23, 52, 4, 5, 24, 6, 42, 28], 3D Gaussians [18, 12, 39], or signed distance fields (SDFs) [46, 45, 36, 49, 50] for images and other optical modalities [7, 14, 30]. While the density fields in NeRF and 3D Gaussians can model arbitrary complex scene geometries, they fundamentally lack well-defined surfaces and normals. These particle-centric representations result in ambiguous surface definitions, which hinders their applicability to PBR. In contrast, SDFs explicitly define surfaces as the zero-level set of a neural network, providing accurate surface and normal estimation. This surface-centric representation makes SDFs inherently more compatible with inverse rendering tasks which rely on PBR and require precise surface definitions.
@@
Figure 2: Overview of our PBR-NeRF architecture for neural inverse rendering. Our two novel physics-based material losses, i.e. $\mathcal{L}_{\text{cons}}$ for conservation of energy and $\mathcal{L}_{\text{spec}}$ for disentangling the diffuse and specular BRDF components, are derived in Sec. 3.2. The complete PBR-NeRF model comprises multiple neural fields, which are optimized in a stage-wise fashion: a standard NeRF+SDF modeling radiance and geometry, a neural incident light field (NeILF) modeling spatially-varying illumination, and a BRDF field modeling materials via the Disney BRDF [11]. Our physics-based losses provide the BRDF field with valuable inductive biases for improved material estimation, which help resolve to a large extent the inherent material-lighting ambiguity in inverse rendering and thus benefit the incident light field as well. Consequently, novel scene views synthesized with our model enjoy state-of-the-art quality.
@@
Neural inverse rendering  predates neural fields [33, 19, 37, 48, 31, 22] and learns to predict 3D geometry, materials, and illumination from 2D images. Decomposed neural fields [35, 38, 47, 27] similarly extend conventional neural fields with inverse rendering or intrinsic decomposition to estimate physical scene properties such as materials and illumination. Compared to NeRF and 3DGS, the above additional material and lighting estimation enables complex downstream tasks such as relighting [43, 29, 21], material editing, and appearance manipulation. The decomposition of reflected radiance into materials and illumination can also address issues such as “baked-in” highlights in predicted colors by properly modeling view-dependent effects. However, inverse rendering introduces significant complexity due to the material-lighting-geometry ambiguity, which is harder to lift than the geometry-radiance ambiguity of standard novel view synthesis. Many approaches simplify this task by assuming known lighting, environment map lighting [53, 54, 25, 8, 9], known geometry [44], or basic Lambertian material models. State-of-the-art inverse rendering approaches jointly model (1) direct and indirect lighting from near and far sources [44, 51, 40], and (2) spatially-varying microfacet BRDFs [34, 44, 51]. For computational efficiency, techniques such as light caching [55, 44, 20, 51, 16, 3] and incident light sampling [44, 51, 3] are crucial to avoid costly recursive light tracing.
@@
NeILF [44] proposed joint incident light and BRDF fields, improving on environment-map-based approaches such as NeRV [34] by handling near-field lights and mixed lighting settings in addition to indirect lighting and occlusions. Building on this, NeILF++ [51] introduced inter-reflectable light fields, combining the NeILF and BRDF fields with an SDF for geometry estimation. This approach uses the SDF’s volume-rendered outgoing radiance to supervise the incident radiance predicted by the NeILF. The reflection loss of [51] harnesses the SDF’s strong novel view synthesis performance to steer the ill-posed inverse rendering, achieving state-of-the-art material estimation among neural inverse rendering works. While largely effective, current material estimation methods including [44, 51] lack informative material priors for disambiguating specular from diffuse properties of scene surfaces. This typically results in incorrectly attributing specular highlights to alleged changes in diffuse albedo, even though such errors do not affect view synthesis results.
@@
Material priors for inverse rendering  are crucial for constraining material estimation and reducing material-lighting ambiguities. BRDF smoothness priors  [44, 51, 16, 3] can discourage abrupt spatial changes in material properties. NeILF [44] and NeILF++ [51] use a Lambertian prior to regularize the Disney BRDF’s metallicness and roughness parameters. However, the Lambertian prior induces a strong assumption of perfectly rough, non-metallic materials that limits BRDF expressiveness. For instance, specular effects often appear in the predicted BRDF’s diffuse lobe due to the Lambertian prior suppressing the specular lobe. We address these limitations by proposing two novel physics-based losses that enhance the physical validity and accuracy of estimated BRDFs without sacrificing expressiveness.
@@ 3.2 Physics-Based Losses
NDF-weighted Specular Loss.  Our second physics-based loss targets the imbalance between the diffuse and specular lobes frequently observed in inverse rendering methods, such as NeILF++ [51], which often assume Lambertian reflection. Real-world materials, however, often violate idealized Lambertian behavior, resulting in “baked-in” specular highlights where the predicted diffuse lobe $f_{d}$ overcompensates for insufficient specular reflection. For these “baked-in” specular highlights, the aggregate BRDF value $f_{r}$ may be accurate for specular directions, but the imbalance between $f_{d}$ and $f_{s}$ causes incorrect diffuse behavior at non-specular angles, degrading material estimation quality.
To correct this imbalance, we propose the NDF-weighted Specular Loss to penalize excessive diffuse reflection in specular regions and we define our loss as
$\mathcal{L}_{\text{spec}}=$
$\frac{1}{|S_{L}|}\sum_{\boldsymbol{\omega}_{i}\in S_{L}}\operatorname{softmax}\left(\frac{\operatorname{detach}\left(D\left(\frac{\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}}{||\boldsymbol{\omega}_{o}+\boldsymbol{\omega}_{i}||}\right)\right)}{T_{\text{spec}}}\right)f_{d},$ (8)
@@
Our NDF-weighted Specular Loss complements the standard RGB rendering loss by pushing the specular lobe $f_{s}$ to compensate for the penalized diffuse lobe $f_{d}$, creating a stronger separation between diffuse and specular reflections. A proper weighting of the opposing RGB rendering loss and other regularization terms ensures that the diffuse lobe is not arbitrarily suppressed by our specular loss. As with $\mathcal{L}_{\text{cons}}$, the specular loss also benefits lighting estimation. An oversized diffuse lobe can lead to underestimated light intensities, and excessive diffuse reflection can also force the estimated incident light to incorrectly reproduce complex, view-dependent specular effects. Together, our two losses $\mathcal{L}_{\text{cons}}$ and $\mathcal{L}_{\text{spec}}$ significantly improve the accuracy of material and lighting estimation, resulting in more realistic scene reconstructions, as we detail in Sec. 4.
@@ 3.3 Joint Optimization
Following NeILF++ [51], we optimize the scene reconstruction in three phases. We briefly summarize the three phases here (See Appendix Sec. D.2 for full details):
@@
$\mathcal{L}_{\text{mat}}=\lambda_{\text{pbr}}\mathcal{L}_{\text{pbr}}+\lambda_{\text{ref}}\mathcal{L}_{\text{ref}}+\lambda_{\text{smth}}\mathcal{L}_{\text{smth}}+\mathcal{L}_{\text{mat,physics}},$ (9)
@@ 4 Experiments
Table 1: Comparison against state-of-the-art methods on the NeILF++ dataset [51]. Material estimation and novel view synthesis are compared across different scenes (City, Studio, and Castel) and types of illumination (Env: global environment map, Mix: mixed lighting with environment map, point, and area light sources) using PSNR and SSIM. †: our reproduced results. N/A: no official reported results.
Predicted Quantity Method Env-City Env-Studio Env-Castel Mix-City Mix-Studio Mix-Castel Mean
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
RGB PhySG [53] 24.82 N/A 25.65 N/A 27.24 N/A 24.38 N/A 24.04 N/A 25.81 N/A 25.32 N/A
SG-ENV [53] 31.01 N/A 29.46 N/A 32.34 N/A 27.20 N/A 25.88 N/A 27.70 N/A 28.93 N/A
NeILF [44] 33.23 92.29 31.12 83.55 36.21 94.13 30.00 86.95 27.56 74.20 30.91 89.83 31.50 86.83
NeILF++  [51] 31.53 91.51 29.77 80.15 33.87 93.26 29.64 88.83 27.50 74.70 30.74 91.07 30.51 86.59
PBR-NeRF (ours) 32.60 92.33 30.96 80.77 34.88 93.66 30.03 89.17 27.84 75.01 31.34 91.39 31.27 87.05
Roughness PhySG [53] 6.62 N/A 11.29 N/A 6.22 N/A 6.27 N/A 6.83 N/A 6.14 N/A 7.23 N/A
SG-ENV [53] 9.61 N/A 17.64 N/A 9.74 N/A 8.77 N/A 12.58 N/A 9.14 N/A 11.25 N/A
NeILF [44] 16.71 86.77 17.72 86.48 18.05 89.56 15.11 81.42 14.54 78.53 15.04 81.28 16.19 84.01
NeILF++  [51] 21.22 91.27 22.01 91.89 19.91 90.63 22.19 91.57 23.43 92.37 22.20 91.61 21.83 91.56
PBR-NeRF (ours) 22.19 92.32 23.45 92.74 22.23 92.07 22.15 92.04 22.63 92.57 22.18 92.10 22.47 92.31
Metallicness PhySG [53] 8.72 N/A 7.97 N/A 8.35 N/A 8.67 N/A 8.95 N/A 8.76 N/A 8.57 N/A
SG-ENV [53] 17.01 N/A 16.40 N/A 16.39 N/A 15.44 N/A 14.25 N/A 14.49 N/A 15.66 N/A
NeILF [44] 19.48 64.43 18.75 92.56 18.21 92.58 21.03 65.53 19.45 89.85 19.60 76.44 19.42 80.23
NeILF++  [51] 19.49 86.50 18.17 92.99 17.41 91.92 20.80 80.48 19.18 68.74 17.98 80.14 18.84 83.46
PBR-NeRF (ours) 21.73 72.66 22.26 73.26 20.88 88.65 22.45 68.40 19.24 63.25 23.15 78.28 21.62 74.08
Albedo PhySG [53] 15.01 N/A 16.96 N/A 16.13 N/A 14.16 N/A 12.43 N/A 14.29 N/A 14.83 N/A
SG-ENV [53] 22.38 N/A 20.74 N/A 22.21 N/A 16.92 N/A 13.16 N/A 16.69 N/A 18.68 N/A
NeILF [44] 17.73 90.32 21.11 90.21 18.24 90.18 17.01 77.14 18.52 81.01 16.32 78.33 18.15 84.53
NeILF++ [51] 18.21 82.62 19.50 79.21 17.73 82.48 16.37 70.13 14.48 64.61 17.43 76.20 17.29 75.87
PBR-NeRF (ours) 19.59 90.60 20.24 88.33 19.88 91.45 20.18 85.53 19.44 77.86 21.14 87.13 20.08 86.82
@@ 4.1 Experimental Setup
Implementation Details.  We implement PBR-NeRF on top of NeILF++ [51] using PyTorch [2] and keep all hyperparameters from NeILF++ the same, except for $|S_{L}|=256$ and a training batch size of 8192 rays. The geometry, material, and joint phases last 5K, 1K, and 30K iterations, respectively. We train with a learning rate of 0.002, which is fixed for the geometry and material phase, but decreases by a factor of 5 every 10K iterations during the joint phase. Training runs on a single NVIDIA A6000, with runtimes ranging from 3 to 7.5 hours. Using grid search, we identify two sets of weights for our physics-based losses: $\lambda_{\text{cons}}=0.01$ and $\lambda_{\text{spec}}=0.5$ for the NeILF++ dataset, and $\lambda_{\text{cons}}=0.01$ and $\lambda_{\text{spec}}=0.01$ for DTU. We weigh $\lambda_{\text{spec}}$ lower than the RGB loss, $\lambda_{\text{pbr}}=1$, so that $\mathcal{L}_{\text{spec}}$ has a small yet meaningful weighting. More hyperparameter details are provided in Appendix Sec. D.3.
Datasets.  We evaluate our method on the NeILF++ [51] and DTU [15] datasets. We use the NeILF++ synthetic dataset to evaluate our material-illumination estimation, assuming known geometry. This dataset provides ground-truth geometry, RGB, albedo, metallicness, and roughness for a scene rendered under 6 static illumination patterns, including three environment maps (Env) and three mixed-lighting scenarios with additional point and area light sources (Mix). Each pattern includes 96 RGB HDR $1600\times 1200$ images, split into 87 training and 9 validation images. Ground-truth geometry and poses are used in training, while ground-truth color, albedo, metallicness, and roughness are only used for evaluation. DTU captures 15 real-world objects in a controlled lab setting. Each scene includes 49 RGB LDR $1600\times 1200$ images, split into 44 training and 5 validation images. We evaluate the predicted geometry by computing the Chamfer Distance with the provided reference point cloud generated by structured light scanners. We only qualitatively evaluate material estimation on DTU because it uses LDR images and does not provide ground-truth materials. We do not evaluate lighting because the DTU dataset is unsuitable for evaluating predicted lighting without HDR images and ground-truth environment maps. Furthermore, the predicting lighting is not meaningful because lighting cannot be reliably estimated from the LDR DTU images due to clipping and unknown tone mapping.
@@
Table 2: Comparison for novel view synthesis against state-of-the-art methods on DTU [15]. Performance is compared across different scenes using the RGB PSNR and Chamfer Distance metrics. †: our reproduced results. $\ddagger$: mean result over 3 runs. *: only 1 of either RGB PSNR or Chamfer Distance results is publicly available.
Method \ DTU scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean±$\pm\sigma$
RGB PSNR NeRFactor [54]* 23.24 21.91 23.33 26.86 22.70 24.71 27.59 22.56 20.45 25.08 26.30 25.14 21.35 26.44 26.53 24.28
PhySG [53]* 17.38 15.11 20.65 18.71 18.89 18.47 18.08 21.98 17.31 20.67 18.75 17.55 21.20 18.78 23.16 19.11
NeILF++ [51$\ddagger$ 25.85 21.93 25.42 27.99 29.76 27.45 26.54 28.37 23.73 28.28 31.56 30.45 26.86 31.32 32.19 27.85$\pm$1.04
PBR-NeRF (ours)$\ddagger$ 26.11 24.49 27.56 28.30 30.13 27.57 27.07 29.89 24.76 28.71 31.83 30.68 27.24 31.64 33.48 28.63 $\pm$ 0.29
Chamfer Distance [mm] VolSDF [46]* 0.810 1.140 0.490 1.250 0.700 0.720 1.290 1.180 1.260 0.700 0.660 1.080 0.420 0.610 0.550 0.857
NeuS [36]* 0.930 1.000 0.430 1.100 0.650 0.570 1.480 1.090 1.370 0.830 0.520 1.200 0.350 0.490 0.540 0.837
NeILF++ [51$\ddagger$ 0.748 3.028 1.285 0.448 1.627 0.640 1.287 1.649 2.697 1.008 0.604 0.745 0.397 0.513 0.824 1.167$\pm$0.68
PBR-NeRF (ours)$\ddagger$ 0.816 1.343 0.486 0.511 1.665 0.688 0.647 1.174 1.416 1.038 0.590 0.676 0.398 0.502 0.468 0.828$\pm$0.09